OpenStudy (anonymous):

Write the equation of a line in slope-intercept form that has a slope of -2 and passes through the point (2, -8).

OpenStudy (anonymous):

OMG it's you again! @texaschic101

OpenStudy (texaschic101):

y = mx + b slope(m) = -2 (2,-8) x = 2 and y = -8 now we will sub... -8 = -2(2) + b (we need to solve for b) -8 = -4 + b -8 + 4 = b -4 = b So the equation of the line you need is : y = -2x - 4 Any questions ?

OpenStudy (texaschic101):

yes it is me....is that a bad thing ??

OpenStudy (anonymous):

No it's not a bad thing, it's just you were here to help me before. I find that kind of funny. I Have literally so many questions I need help with I don't understand slope at all.

OpenStudy (texaschic101):

if you have any more problem that you need help with, ask away. I will help with what I can :)

OpenStudy (anonymous):

Thanks so much. Uhm.. Write the equation of a line parallel to the line 2x + 3y = 5 and passes through the point (9, -3)

OpenStudy (texaschic101):

first we need to find the slope. To do that, we need to put the equation in y = mx + b form and m is your slope. 2x + 3y = 5 (subtract 2x from both sides) 3y = -2x + 5 (divide both sides by 3) y = -2/3x + 5/3 your slope is -2/3. We are looking for a parallel line, so the slope will stay the same. now we use y = mx + b slope(m) = -2/3 (9,-3) x = 9 and y = -3 now we sub.. y = mx + b -3 = -2/3(9) + b (we are solving for b(the y intercept) -3 = -18/3 + b -3 = - 6 + b (add 6 to both sides) -3 + 6 = b 3 = b your parallel line is : y = -2/3x + 3

OpenStudy (anonymous):

o__o Thank you. Though tbh I still don't understand how to do these. It may be that I get stuck on putting Y=mx+b into the equation.

OpenStudy (texaschic101):

there is more then 1 way to do this, although you still have to find the slope using the y = mx + b form. Once you have the slope, you can either use y = mx + b and solve for b, or you can use y - y1 = m(x - x1). Either way you do it, you will still get the same answer. It just takes practice

OpenStudy (anonymous):

Write an equation of a line in point-slope form that has a slope of -2 and passes through (5, -1).

OpenStudy (texaschic101):

are you having problems putting the equation in y = mx + b form to find the slope ?

OpenStudy (texaschic101):

y = mx + b slope(m) = -2 (5,-1) x = 5 and y = -1 can you try....just sub in -1 for y, -2 for m, and 5 for x....then solve for b.

OpenStudy (texaschic101):

wait.......It is different on this one...sorry..hold on

OpenStudy (anonymous):

Yes I am. Sadly I've always had this problem. Okay will do.

OpenStudy (texaschic101):

point slope form is : y - y1 = m(x - x1) slope(m) = -2 (5,-1) x1 = 5 and y1 = -1 now we sub... y - (-1) = -2(x - 5) y + 1 = -2(x - 5) <-- your answer when it asks for point slope form, it has to be done this way. You do not have to solve the formula, in fact, you leave the parenthesis in. Does this make sense to you ?

OpenStudy (anonymous):

I am going to have to get off of here until later. I really appreciate it. If you're on later. I have several more questions I could you explaining on thanks. Eh maybe a bit. D: Sorry.

OpenStudy (anonymous):

Rewrite the equation of a line in point-slope form to slope-intercept form: y minus 3 equals one-half times the quantity x minus 8

OpenStudy (anonymous):

So would ^^ be. . y equals one-half x minus 5

OpenStudy (texaschic101):

y - 3 = 1/2x - 5 (add 3 to both sides) y = 1/2x - 5 + 3 y = 1/2x - 2 you can't forget about the 3, you just have to move it to the other side.

OpenStudy (texaschic101):

@Maryok ....^^^ :)

OpenStudy (anonymous):

That isn't one of the choices. D:

OpenStudy (anonymous):

Or would it be - 5?

OpenStudy (texaschic101):

my bad...I read the question wrong...let me try again..give me a sec

OpenStudy (anonymous):

Thank you!!

OpenStudy (texaschic101):

y - 3 = 1/2(x - 8) -- distribute through the parenthesis y - 3 = 1/2x - 4 -- add 3 to both sides y = 1/2x - 4 + 3 -- combine like terms y = 1/2x - 1 understand ??

OpenStudy (anonymous):

Lmfao no I don't.

OpenStudy (texaschic101):

really ? What are you having problems with ?

OpenStudy (anonymous):

Could you help me on 3 more though? I'm not to sure I guess it's just the whole concept.

OpenStudy (texaschic101):

yes...let me see what you got

OpenStudy (anonymous):

Determine if lines that pass through the given points are parallel, perpendicular or neither. Line A: (5, 4) and (7, 5) Line B: (4, -5) and (8, -3) A. Parallel B. Perpendicular C. Neither D. Cannot be determined

OpenStudy (anonymous):

Do I just graph these and see?

OpenStudy (texaschic101):

we need to find the slopes of these two lines by using the slope formula. If the slopes are the same, it is parallel. If the slopes are negative reciprocals of each other, then they are perpendicular. I suppose you can graph it and find out or we can solve for the slopes....what do you want to do ? Just to let you know, I am not that good at graphing

OpenStudy (anonymous):

Solving will do. Let me see if I could possibly come up with anything.__.

OpenStudy (anonymous):

and...I've come up with nothing. I don't know.

OpenStudy (texaschic101):

slope formula is : slope(m) = (y2 - y1) / (x2 - x1) first line... (5,4) x1 = 5 and y1 = 4 (7,5) x2 = 7 and y2 = 5 now we sub... slope(m) = (y2 - y1) / (x2 - x1) m = (5 - 4) / (7 - 5) m = 1/2 <-- this is the slope for Line A. second line... (4,-5) x1 = 4 and y1 = -5 (8,-3) x2 = 8 and y2 = -3 time to sub again slope(m) = (y2 - y1) / (x2 - x1) m = (-3-5) / (8 - 4) m = -8/4 m = -2 slope of A = 1/2 slope of B = -2 what we have here is a negative reciprocal....to find the negative reciprocal of a number, you " flip " the slope and change the sign. The negative reciprocal of 1/2 - -2/1 or just -2 (you see how I flipped the slope and changed the sign.) That makes these line perpendicular.

OpenStudy (texaschic101):

oops..typo...the negative reciprocal of 1/2 = -2/1 or just -2

OpenStudy (anonymous):

I so don't understand how you get this. Also thanks for being so patient, and explaining everything.

OpenStudy (texaschic101):

I wish I could just sit down with you on one one. I could explain this stuff to you and you could understand it better.

OpenStudy (texaschic101):

oops...one on one

OpenStudy (texaschic101):

what you need to learn is : slope formula -- slope(m) = (y2 - y1) / (x2 - x1) point slope formula -- y - y1 = m(x - x1) slope intercept formula -- y = mx + b

OpenStudy (anonymous):

I know as do I. See I've learned everything, and have always had someone like you said sit down with me one on one, and show me no matter how long it took. It's just that I'm doing Online school this year, and don't have that...and I just don't get it unless it's explained.

OpenStudy (texaschic101):

also, parallel lines have the same slope and perpendicular lines have negative reciprocal slopes

OpenStudy (anonymous):

Write the equation of a line in point-slope form that has a slope of one-third and passes through the point (-6, 0).

OpenStudy (texaschic101):

I am sure there is websites you can go over, but I still don't think it is the same as working one on one with a person....let me check out your next problem....

OpenStudy (anonymous):

and there are. I just can't explain it to myself I have to have it explained.

OpenStudy (texaschic101):

point slope form-- y - y1 = m(x - x1) slope(m) = 1/3 (-6,0) x1 = -6 and y1 = 0 now we just sub.. y - 0 = 1/3(x - (-6) or it could also be written as : y - 0 = 1/3(x + 6)

OpenStudy (texaschic101):

can you kinda understand what I did ?

OpenStudy (anonymous):

Maybe a bit. I just don't...see it.

OpenStudy (texaschic101):

do you understand that the set of points are always in (x,y) form ?

OpenStudy (texaschic101):

and in y = mx + b, the m is the slope and the b is the y intercept

OpenStudy (anonymous):

Hahah I don't.

OpenStudy (anonymous):

Yes I understand that, and to find the slope it is basically Rise/Run

OpenStudy (texaschic101):

true...rise/run. But to find the slope without looking at a graph and just getting two sets of points, you have to use the slope formula.

OpenStudy (texaschic101):

personally, I do not see that I will be using this stuff in the future, but for school, you still have to learn it

OpenStudy (texaschic101):

you got anymore problems ?

OpenStudy (anonymous):

Yes sadly. What is the x-intercept of the graph: 2x + 4y = -10

OpenStudy (texaschic101):

to find the x intercept, just sub in 0 for y 2x + 4y = -10 2x + 4(0) = -10 2x + 0 = -10 2x = -10 x = -10/2 x = -5 and if you wanted to find the y intercept, you could sub in 0 for x

OpenStudy (anonymous):

A. (0, -5) B. (-5, 0) C. the ordered pair 0, negative five-halves D. the ordered pair negative 5, negative five-halves

OpenStudy (texaschic101):

ok...x intercept is (-5,0) because your set of points is (x,y)

OpenStudy (texaschic101):

and x = -5 and y = 0

OpenStudy (texaschic101):

anything else ?

OpenStudy (anonymous):

Yes I'm trying to work it real quick.

OpenStudy (texaschic101):

I understand...take your time..I am in no hurry

OpenStudy (anonymous):

What is the y-intercept of the graph: 3x – 2y = 12 A. (4, 0) B. (4, -6) C. (0, 6) D. (0, -6)

OpenStudy (anonymous):

So this would be A?

OpenStudy (texaschic101):

to find the y intercept, sub in 0 for x 3x - 2y = 12 3(0) - 2y = 12 0 - 2y = 12 -2y = 12 y = -12/2 y = -6......which is (0,-6) answer D

OpenStudy (texaschic101):

remember, to find the y intercept, sub in 0 for x and to find the x intercept, sub in 0 for y. Then put your answer accordingly (x,y)

OpenStudy (anonymous):

Right sorry!

OpenStudy (texaschic101):

no need to apologize...you did nothing wrong

OpenStudy (texaschic101):

your just learning

OpenStudy (texaschic101):

believe me, if I had $1 for every mistake I made, I would be rich...lol

OpenStudy (anonymous):

Thank you. Hahaha. As would I. If you don't mind me asking are you a teacher or?

OpenStudy (texaschic101):

no...I am not a teacher....I would probably not be a good one, I would probably yell to much..lol

OpenStudy (anonymous):

Well. The only reason I asked was because you're quite patient, or so it seems.

OpenStudy (anonymous):

How would you graph..Graph the line y = -3x -2

OpenStudy (texaschic101):

It seems....on the computer, I can be patient to a point...in person, 8 hours a day with a classroom of students would drive me crazy

OpenStudy (anonymous):

I'd expect it would. Don't understand how they do it.

OpenStudy (texaschic101):

graph...y = -3x - 2 In y = mx + b form, the m is the slope and the y intercept is b. Therefore, plot your point (0,-2)-- the y intercept,which is on the y line. To find the x intercept, sub in 0 for y y = -3x - 2 0 = -3x - 2 3x = -2 x = -2/3 (plot your point (-2/3,0) on the x line. and you can draw your line connecting the two points. And if it helps, the slope = -3, which is the same as -3/1.....so from point (0,-2) come down 3(because it is negative) and to the right 1(because it is positive), then down 3 and to the right 1, and so on, until you cross the x axis at (-2/3,0) It looks like this..... http://www.quickmath.com/webMathematica3/quickmath/graphs/equations/basic.jsp#c=graph_stepsgraphequation&v1=y+%3D+-3x+-+2 does that make sense ?

OpenStudy (texaschic101):

since the y intercept is negative, you can do the opposite of the slope....up 3 and to the left 1 and up 3 and to the left 1...and you have your line

OpenStudy (texaschic101):

I am not the best at explaining graphing...sorry about that...I probably confused the hell out of you

OpenStudy (anonymous):

Nah you didn't you're actually helping me out A LOT...and it means A LOT. LOL

OpenStudy (texaschic101):

glad I could help....any more questions ?

OpenStudy (anonymous):

No I believe that is all I'm probably going to stop here, and move on to English. Thanks again. (,:

OpenStudy (texaschic101):

no problem....good luck learning this stuff. Come back to this post and review it if you need to. Good luck at english too.....I hate english class...lol

OpenStudy (anonymous):

Hahah Thank you. I shall. Tell me about it. Trust me I will probably be back tomorrow with hundreds of questions.

OpenStudy (texaschic101):

I will probably be on here sometime tomorrow..I can help

OpenStudy (anonymous):

That would be wonderful. Hah